OpenStudy (anonymous):

In a regular hexagonal pyramid, each lateral face is inclined at an angle of 60 degrees with the base. if the are of the base is 200 sqrt of 3 ft^2. Find the volume of the pyramid.

OpenStudy (shinalcantara):

Area of a regular polygon is defined by the formula \[A = \frac{ 1 }{ 2 }nR^2\sin \theta\] where n is the number of sides R is the radius of the circle that can be circumscribing the polygon theta is 360/n \[A = 200\sqrt{3}\] \[n=6\] \[\theta = \frac{ 360^o }{ 6 } \rightarrow 60^o\] \[200\sqrt{3} = \frac{ 1 }{ 2 }6R^2\sin60^o\] \[R^2 = \frac{ 200\sqrt{3(2)} }{ 6\sin60^o }\] \[R^2 = \frac{ 400 }{ 3 }\] \[R = \frac{ 20\sqrt{3} }{ 3 }\]

OpenStudy (shinalcantara):

internet connection's annoying -_-

OpenStudy (shinalcantara):

The main purpose of getting R is to find 'x' and eventually 'h' or the altitude of the pyramid. Hence the formula for the volume of pyramid is\[V = \frac{ 1 }{ 3 } Bh\] where B is the area of the base and h is the altitude. Area is given, so all we have to know is the height.

OpenStudy (shinalcantara):

find 'x' using trigonometry \[\cos30^o = \frac{ x }{ R }\] \[x = Rcos30^o\] \[x = \frac{ 20\sqrt{3} }{ 3 }\cos30^o\] \[x = 10\] Since the lateral faces are inclined at 60 degrees then the height can be solved \[\tan60^o = \frac{ h }{ x }\] \[h = xtan60^o\] \[h = 10\tan60^o\] \[h = 10\sqrt3\] Volume then would be: \[V = \frac{ 1 }{ 3 }Bh\] \[V = \frac{ 1 }{ 3 } (200\sqrt{3}) (10\sqrt{3})\] \[V = 2,000 ft^3\]